# Stopping rules for Monte Carlo methods of martingale difference type

### v1

@@ B.1 Stopping rules in asymptotic regimes
We have justified our formulations in an asymptotic manner along the way (Propositions 4.3 and 4.5) in a similar spirit to the existing framework in the asymptotic regime [12]. A key assumption can be rephrased in our context that the batch variances are convergent. That is, given that a limiting unit-time variance exits ($\lim_{t\to+\infty}v^{2}_{0}(t)(=:c)$), the strong consistency holds ($v_{1}^{2}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}c$, or $v_{2}^{2}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}c$), or the functional weak law of large numbers holds ($\{s^{2}(\lfloor t/\epsilon\rfloor):\,t\in(0,+\infty)\}\stackrel{{\scriptstyle\mathcal{L}}}{{\to}}c$, or $\{\sigma^{2}(\lfloor t/\epsilon\rfloor):\,t\in(0,+\infty)\}\stackrel{{\scriptstyle\mathcal{L}}}{{\to}}c$ in $\mathbb{D}((0,+\infty);\mathbb{R})$). Clearly, all those conditions are stronger than ours (Assumption 4.1 and the additional conditions in Proposition 4.5). Another key condition is the existence of a limiting continuous process of the scaled-centered estimation process weakly in $\mathbb{D}((0,+\infty);\mathbb{R}))$ in conjunction with the limiting unit-time variance. To verify this condition in our context, consider the (non-batched) empirical mean $\mu_{n}:=n^{-1}\sum_{k=1}^{n}X_{k}$, instead of the batched one (2.2), so that $\mu_{n}\to\mu$ almost surely by Theorem 2.4. For $\epsilon\in(0,+\infty)$, define stochastic process $\{\mathcal{Y}_{\epsilon}(t):\,t\in(0,+\infty)\}$ by
@@
for all $0\leq t_{1}\leq t_{2}\leq t_{3}\leq t_{4}$. To further show that the limiting process is a Brownian motion, let $\{\epsilon_{n}\}_{n\in\mathbb{N}}$ and $\{h_{n}\}_{n\in\mathbb{N}}$ be sequences of positive constants with $\epsilon_{n}\to 0$ and $h_{n}\to 0$ as $n\to+\infty$ and let $\tau_{n}$ be a stopping time with respect to the stochastic process $\{\mathcal{X}_{\epsilon_{n}}(t):\,t\geq 0\}$ for all $n\in\mathbb{N}$, that is, $\{\tau_{n}\leq t\}\in\sigma(\{\mathcal{X}_{\epsilon_{n}}(s):\,s\in[0,t]\})$. Due to the martingale central limit theorems (Theorems 4.2, 4.4 and 4.6), the limiting process is Gaussian and $\{\mathcal{X}_{\epsilon_{n}}(t):\,t\geq 0\}$ converges weakly in $\mathbb{D}([0,+\infty);\mathbb{R})$ to a centered Brownian motion, due to tightness $\mathbb{P}_{0}(|\mathcal{X}_{\epsilon_{n}}(\tau_{n}+h_{n})-\mathcal{X}_{\epsilon_{n}}(\tau_{n})|>m)\lesssim m^{-2}h_{n}\lim_{t\to+\infty}v^{2}_{0}(t)\to 0$ as $n\to+\infty$ for all $m>0$. (See, for instance, [20, Chapter 23].)
@@ B.2 Relative-precision stopping rules
where we have employed the key fact that the resampled batch $\{X_{k}^{\star}\}_{k\in M(\tau)}$ is conditionally independent of the stopping batch $\tau$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau)}$ for the martingale central limit theorem (Theorem 4.2). The crucial issue here lies in how to treat the unknown mean $\mu$ in the rightmost term, which is exactly what one is looking for. On the one hand, it cannot be replaced with the original empirical mean $\mu(\tau)$ because it is not clear whether $\mu(\tau)$ approximates $\mu$. We stress, on the other hand, that $\mu$ there cannot be replaced with the resampled empirical mean $\mu_{\star}(\tau)$ either despite it approximates $\mu$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau)}$, since then the stopping batch $\tau$ would be determined after generating the resampled one $\{X_{k}^{\star}\}_{k\in M(\tau)}$ in reverse order.
Before closing this section, let us note, as already outlined along Assumption 2.1, that a relative-precision stopping rule was developed in [9] upon a similar assumption to (2.1), but along with a boundedness condition on the underlying randomness (specifically, $X_{n}\in[0,1]$, $a.s.$), which plays an essential role for deriving bounds for the error probability upon which the stopping rule is built.

### v2

@@ B.1 Stopping rules in asymptotic regimes
We have justified our formulations in an asymptotic manner along the way (Propositions 4.3 and 4.5) in a similar spirit to the existing framework in the asymptotic regime [PW1992]. A key assumption can be rephrased in our context that the batch variances are convergent. That is, given that a limiting unit-time variance exits ($\lim_{t\to+\infty}v^{2}_{0}(t)(=:c)$), the strong consistency holds ($v_{1}^{2}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}c$, or $v_{2}^{2}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}c$), or the functional weak law of large numbers holds ($\{s^{2}(\lfloor t/\epsilon\rfloor):\,t\in(0,+\infty)\}\stackrel{{\scriptstyle\mathcal{L}}}{{\to}}c$, or $\{\sigma^{2}(\lfloor t/\epsilon\rfloor):\,t\in(0,+\infty)\}\stackrel{{\scriptstyle\mathcal{L}}}{{\to}}c$ in $\mathbb{D}((0,+\infty);\mathbb{R})$). Clearly, all those conditions are stronger than ours (Assumption 4.1 and the additional conditions in Proposition 4.5). Another key condition is the existence of a limiting continuous process of the scaled-centered estimation process weakly in $\mathbb{D}((0,+\infty);\mathbb{R}))$ in conjunction with the limiting unit-time variance. To verify this condition in our context, consider the (non-batched) empirical mean $\mu_{n}:=n^{-1}\sum_{k=1}^{n}X_{k}$, instead of the batched one (2.2), so that $\mu_{n}\to\mu$ almost surely by Theorem 2.4. For $\epsilon\in(0,+\infty)$, define stochastic process $\{\mathcal{Y}_{\epsilon}(t):\,t\in(0,+\infty)\}$ by
@@
for all $0\leq t_{1}\leq t_{2}\leq t_{3}\leq t_{4}$. To further show that the limiting process is a Brownian motion, let $\{\epsilon_{n}\}_{n\in\mathbb{N}}$ and $\{h_{n}\}_{n\in\mathbb{N}}$ be sequences of positive constants with $\epsilon_{n}\to 0$ and $h_{n}\to 0$ as $n\to+\infty$ and let $\tau_{n}$ be a stopping time with respect to the stochastic process $\{\mathcal{X}_{\epsilon_{n}}(t):\,t\geq 0\}$ for all $n\in\mathbb{N}$, that is, $\{\tau_{n}\leq t\}\in\sigma(\{\mathcal{X}_{\epsilon_{n}}(s):\,s\in[0,t]\})$. Due to the martingale central limit theorems (Theorems 4.2, 4.4 and 4.6), the limiting process is Gaussian and $\{\mathcal{X}_{\epsilon_{n}}(t):\,t\geq 0\}$ converges weakly in $\mathbb{D}([0,+\infty);\mathbb{R})$ to a centered Brownian motion, due to tightness $\mathbb{P}_{0}(|\mathcal{X}_{\epsilon_{n}}(\tau_{n}+h_{n})-\mathcal{X}_{\epsilon_{n}}(\tau_{n})|>m)\lesssim m^{-2}h_{n}\lim_{t\to+\infty}v^{2}_{0}(t)\to 0$ as $n\to+\infty$ for all $m>0$. (See, for instance, [kallenberg, Chapter 23].)
@@ B.2 Relative-precision stopping rules
where we have employed the key fact that the regenerated batch $\{X_{k}^{\star}\}_{k\in M(\tau)}$ is conditionally independent of the stopping batch $\tau$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau)}$ for the martingale central limit theorem (Theorem 4.2). The crucial issue here lies in how to treat the unknown mean $\mu$ in the rightmost term, which is exactly what one is looking for. On the one hand, it cannot be replaced with the original empirical mean $\mu(\tau)$ because it is not clear whether $\mu(\tau)$ approximates $\mu$. We stress, on the other hand, that $\mu$ there cannot be replaced with the regenerated empirical mean $\mu_{\star}(\tau)$ either despite it approximates $\mu$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau)}$, since then the stopping batch $\tau$ would be determined after generating the regenerated one $\{X_{k}^{\star}\}_{k\in M(\tau)}$ in reverse order.
Before closing this section, let us note, as already outlined along Assumption 2.1, that a relative-precision stopping rule was developed in [LWP2013] upon a similar assumption to (2.1), but along with a boundedness condition on the underlying randomness (specifically, $X_{n}\in[0,1]$, $a.s.$), which plays an essential role for deriving bounds for the error probability upon which the stopping rule is built.